# Data Augmentation for Time-Series Classification: a Comprehensive Survey

### 20 Mar 2024 v3

@@ Abstract
Data Augmentation (DA) has emerged as an indispensable strategy in Time Series Classification (TSC), primarily due to its capacity to amplify training samples, thereby bolstering model robustness, diversifying datasets, and curtailing overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible, user-oriented tools. In light of these challenges, this study embarks on an exhaustive dissection of DA methodologies within the TSC realm. Our initial approach involved an extensive literature review spanning a decade, revealing that contemporary surveys scarcely capture the breadth of advancements in DA for TSC, prompting us to meticulously analyze over 100 scholarly articles to distill more than 60 unique DA techniques. This rigorous analysis precipitated the formulation of a novel taxonomy, purpose-built for the intricacies of DA in TSC, categorizing techniques into five principal echelons: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. Our taxonomy promises to serve as a robust navigational aid for scholars, offering clarity and direction in method selection. Addressing the conspicuous absence of holistic evaluations for prevalent DA techniques, we executed an all-encompassing empirical assessment, wherein upwards of 15 DA strategies were subjected to scrutiny across 8 UCR time-series datasets, employing ResNet and a multi-faceted evaluation paradigm encompassing Accuracy, Method Ranking, and Residual Analysis, yielding a benchmark accuracy of 88.94 ± 11.83%. Our investigation underscored the inconsistent efficacies of DA techniques, with certain strategies such as Permutation (boasting an average rank of 4.125 and an accuracy pinnacle of 89.20 ± 11.95%, and outperforming in 2 datasets) significantly propelling model performance, whereas others, like Rotation (84.75 ± 12.92%), inadvertently compromised it. Furthermore, we identified that the inherent attributes of datasets exert a profound impact on the success of DA, leading us to provide precise, pragmatic recommendations, derived from our empirical insights, to assist practitioners in aligning their choice of DA techniques with specific dataset nuances.
@@ I-B Objectives and Contributions
• Comprehensive Survey & Novel Taxonomy: Endeavoring to impose order and coherence upon the diverse techniques permeating this field, we unveil a pioneering taxonomy tailored explicitly for DA in TSC. Our classification schema compartmentalizes methods into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, thereby facilitating a methodical and streamlined approach to DA for TSC study. This structured paradigm is instrumental in the comparative analysis and judicious selection of techniques, calibrated to the nuances of specific tasks.
• Detailed Evaluation: In a move transcending the conventional survey and taxonomy, this study makes a seminal contribution by undertaking a rigorous evaluation of more than fifteen prominent DA for TSC methodologies. These methods were subjected to a battery of tests employing a proven neural network model, establishing a robust and reliable performance benchmark. Our empirical scrutiny provides an all-encompassing perspective on the relative merits and demerits of these techniques, yielding critical insights of immense value to both the scholarly and practical realms.
• Unified Library: Complementing these efforts, we have orchestrated a Python library, encapsulating the aforementioned time series data augmentation algorithms, designed for effortless implementation. This initiative stems from the recognition of the existing tools’ non-intuitive nature, necessitating a more user-friendly alternative.
@@
The rest of this article is organized as follows (Figure 1). In Section 2, we introduce the literature review of DA for TSC. Section 3 presents the survey methodology of our works and discussion of the result. In Section 4, we present the taxonomy for DA techniques, which compartmentalizes the techniques into five principal categories. In Section 5, we introduce the experimental study setup in detail and raise our research questions. After that, the research questions are answered in Section 6 and we discuss the limitations and threats to validity in Section 7. Finally, conclusions are drawn in Section 8.
@@ II-A Relevant Survey & Study
While these studies offer valuable insights into data augmentation techniques, there is a conspicuous absence of a holistic view encompassing recent advancements and diverse application fields. This paper seeks to bridge this scholarly gap, providing an exhaustive analysis of the current state of time series data augmentation, and serving as a compass for future research trajectories in this burgeoning field.
@@ VI-C1 Dataset Characteristics Analysis
In addressing RQ2 regarding the influence of dataset characteristics on augmentation effectiveness and user recommendations: Dataset attributes critically dictate the appropriateness and success of data augmentation strategies. The time series length significantly influences augmentation selection, as seen with ScreenType and ECG200, where Time Warping and Window Slicing’s effectiveness varied markedly. The quantity of training data is another decisive factor, as evidenced by Permutation’s varying impact on CBF and FordB. Additionally, the dataset’s nature and its sensitivity to noise influence augmentation compatibility, as demonstrated with Strawberry and Jitter. Furthermore, the class diversity within datasets like ECG5000 necessitates augmentations that bolster class distinctions.

### 09 Apr 2024 v4

@@ Abstract
Data Augmentation (DA) has emerged as an indispensable strategy in time series classification (TSC), mainly due to its ability to amplify training samples, thus bolstering model robustness, diversifying datasets, and curtailing overfitting. However, the current landscape of DA in TSC is plagued with fragmented literature reviews, nebulous methodological taxonomies, inadequate evaluative measures, and a dearth of accessible and user-oriented tools. In light of these challenges, this study embarks on an exhaustive dissection of DA methodologies within the TSC realm. Our initial approach involved an extensive literature review that spanned a decade, revealing that contemporary surveys scarcely capture the breadth of advancements in DA for TSC, requiring us to meticulously analyze more than 100 scholarly articles to distill more than 60 unique DA techniques. This rigorous analysis precipitated the formulation of a novel taxonomy, designed for the intricacies of DA in TSC, categorizing techniques into five principal echelons: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation. Our taxonomy promises to serve as a robust navigational aid for scholars, offering clarity and direction in method selection. Addressing the conspicuous absence of holistic evaluations for prevalent DA techniques, we carried out an all-encompassing empirical assessment, in which more than 15 DA strategies were subjected to scrutiny across 8 UCR time-series datasets, employing ResNet and a multifaceted evaluation paradigm that includes accuracy, method classification and residual analysis, resulting in a benchmark accuracy of 88.94 ± 11.83%. Our investigation underscored the inconsistent efficacies of DA techniques, with certain strategies such as Permutation (boasting an average rank of 4.125 and an accuracy pinnacle of 89.20 ± 11.95%, and outperforming in 2 datasets) significantly propelling model performance, whereas others, like Rotation (84.75 ± 12.92%), inadvertently compromised it. Furthermore, we identified that the inherent attributes of datasets exert a profound impact on the success of DA, leading us to provide precise and pragmatic recommendations, derived from our empirical insights, to help practitioners align their choice of DA techniques with specific dataset nuances. To improve practical use, we’ve consolidated most of these methods into a unified Python package, whose user-friendly interface facilitates experimenting with various augmentation techniques, offering practitioners and researchers a more convenient tool for innovation than currently available options.
@@ I-B Objectives and Contributions
1. Comprehensive Survey & Novel Taxonomy: Endeavoring to impose order and coherence upon the diverse techniques permeating this field, we unveil a pioneering taxonomy tailored explicitly for DA in TSC. Our classification schema compartmentalizes methods into quintessential categories: Transformation-Based, Pattern-Based, Generative, Decomposition-Based, and Automated Data Augmentation, thereby facilitating a methodical and streamlined approach to DA for TSC study. This structured paradigm is instrumental in the comparative analysis and judicious selection of techniques, calibrated to the nuances of specific tasks.
2. Detailed Evaluation: In a move transcending the conventional survey and taxonomy, this study makes a seminal contribution by undertaking a rigorous evaluation of more than fifteen prominent DA for TSC methodologies. These methods were subjected to a battery of tests employing a proven neural network model, establishing a robust and reliable performance benchmark. Our empirical scrutiny provides an all-encompassing perspective on the relative merits and demerits of these techniques, yielding critical insights of immense value to both the scholarly and practical realms.
3. Unified Library: Complementing these efforts, we have orchestrated a Python library, encapsulating the aforementioned time series data augmentation algorithms, designed for effortless implementation. This initiative stems from the recognition of the existing tools’ non-intuitive nature, necessitating a more user-friendly alternative.
@@
The structure of this article is meticulously organized to facilitate a comprehensive understanding of our investigation, as depicted in Figure 1. Section 2 embarks on a thorough literature review, delineating the state of the art in Domain Adaptation (DA) for Time Series Classification (TSC), establishing the foundation upon which our study is predicated. In Section 3, we elaborate on the methodological framework adopted for our survey, including the criteria for selection and analysis of relevant works, and engage in a detailed discussion of the findings derived therefrom. Section 4 introduces a novel taxonomy for DA techniques, methodically categorizing these methodologies into five distinct principal categories, thereby offering a structured lens through which the current landscape can be viewed. Section 5 is dedicated to the elucidation of our experimental design, meticulously detailing the setup employed in our empirical study and articulating the research questions poised at the outset. Subsequently, Section 6 addresses these research questions, presenting a critical analysis of the experimental outcomes and drawing insights pertinent to the field of DA for TSC. Section 7 ventures into an exploration of the limitations inherent to our study and the potential threats to the validity of our findings, fostering a transparent discourse on the scope and applicability of our conclusions. The article culminates in Section 8, where we synthesize our contributions to the domain and outline directions for future research.
@@ II-A Relevant Survey & Study
Although these studies offer valuable insights into data augmentation techniques, there is a conspicuous absence of a holistic view that encompasses recent advances and diverse application fields. Several data augmentation techniques referenced in the literature are either antiquated or no longer align with contemporary research and practice standards, thereby failing to offer a user-friendly and accessible toolset for researchers and practitioners. This article endeavors to address this gap in the scholarly discourse by conducting a comprehensive analysis of the current landscape of data augmentation within the domain of time-series classification. It aims to serve as a pivotal reference point, guiding future research directions and fostering innovation in this rapidly evolving field.
@@ VI-B4 Detailed Analysis of Top-Performing and Underperforming Data Augmentation Techniques
In response to RQ1 regarding the universality and optimality of data augmentation techniques for time-series classification:
@@ VI-C1 Dataset Characteristics Analysis
In addressing RQ2 regarding the influence of dataset characteristics on augmentation effectiveness and user recommendations:
Dataset attributes critically dictate the appropriateness and success of data augmentation strategies. The time series length significantly influences augmentation selection, as seen with ScreenType and ECG200, where Time Warping and Window Slicing’s effectiveness varied markedly. The quantity of training data is another decisive factor, as evidenced by Permutation’s varying impact on CBF and FordB. Additionally, the dataset’s nature and its sensitivity to noise influence augmentation compatibility, as demonstrated with Strawberry and Jitter. Furthermore, the class diversity within datasets like ECG5000 necessitates augmentations that bolster class distinctions.